In an arithmetic series the sum of T2 + T5 is 18. The 6th Term is greater than the third by 9.
Find the 1st Term and common difference.


Answer by richwmiller(17219)

an = a + (n - 1)d
a2 = a + d
a5 = a + (4)d
18= a + d+a + 4d=2a+5d
a6 = a + 5d
a9 = a + 8d
a + 5d-(a + 8d)=9
-3d=9
d=-3 common difference
18= 2a+5d
18= 2a-15
33=2a
a1=16.5 first term
check
a2=13.5
a5=16.5+4*(-3)=4.5
a6=16.5-15=1.5
a9=16.5-24=-7.5
1.5-(-7.5)=9
ok